An experiment requires a solution that is 80% methyl alcohol by volume. What volume of methyl alcohol should be added to 200 mL of water to make this solution?
1
Expert's answer
2019-11-21T06:26:06-0500

The percentage of CH3OH by volume is defined as:


"\phi_{MeOH}=\frac{V_{MeOH}}{V_{MeOH}+V_{H_2O}}*100\%;"


We have to modify this equation:


"V_{MeOH}*\phi_{MeOH}+\phi_{MeOH}*V_{H_2O}=100\%*V_{MeOH};"


"V_{MeOH}(100\%-\phi_{MeOH})=\phi_{MeOH}*V_{H_2O};"


"V_{MeOH}=\frac{\phi_{MeOH}*V_{H_2O}}{100\%-\phi_{MeOH}}=\frac{80\%*200mL}{100\%-80\%}=800mL."


Thus, 800 mL of pure methyl alcohol should be added to 200 mL of water.